An experienced accountant can prepare a tax return in 13 hours. A novice accountant can do the job in 20 hours. How long will it take them to do the job working together?
Answer by stanbon(75887)

An experienced accountant can prepare a tax return in 13 hours. A novice accountant can do the job in 20 hours. How long will it take them to do the job working together?
-----------
Exp account DATA:
time = 13 hr/job ; rate= 1/13 job/hr
------------
Novice account DATA:
time = 20 hr/job ; rate= 1/20 job/hr
---------------
Together DATA:
time = x hr/job ; rate = 1/x job/hr
-----------------
EQUATION:
rate + rate = rate together
1/13 + 1/20 = 1/x
33/260 = 1/x
260/33 = x
x= 7.88 hrs (time required working together)
=============
Cheers,
Stan H.